# Correction to Hawking radiation in non-singular gravitational collapse

### v2

@@ IV Spontaneous emission as tunneling
    In this formalism, the propagator for a relativistic particle is given by $\braket{x|x^{\prime}}=\int d\lambda_{0}\leavevmode\nobreak\ e^{-im^{2}\lambda_{0}}F(\lambda_{0},x,x^{\prime})$, where $F(\lambda_{0},x,x^{\prime})$ is the amplitude of propagation from $x^{\prime}$ to $x$ in parameter time $\lambda_{0}$.
@@
    The paths for which the Hamilton-Jacobi equation holds are null. Tunneling involves traversing classically forbidden regions using complex paths (Refs. [55, Chapter VII] and [56]).
@@
Before closing this section, a historical note is in order. The Hamilton-Jacobi method, encapsulated in (IV.3) and the picture of particle emission by a black hole as spontaneous pair creation, was first employed in the study of Hawking radiation by Parikh and Wilczek [48], and has been considerably refined since then (see Ref. [35] for a comprehensive review). However, the typical starting point in all these works on black holes is the formula (IV.3), its origin left unexplained. In this and the preceding section, besides obtaining the probability of pair creation by non-singular gravitational collapse, an attempt has been made to provide the link between (IV.3) and the understanding of particle emission by black holes as spontaneous pair creation.
@@ V Comparison with the classical result
$\sigma\sim\exp\left(-\frac{2\pi\omega_{\text{NEH}}}{\kappa_{\text{NEH}}}\right)\leavevmode\nobreak\ .$ (V.1)
@@
$\sigma\sim\exp\left(-8\pi M_{\text{ADM}}\leavevmode\nobreak\ \omega_{\infty}\right)$ (V.2)
By invoking the principle of detailed balance [35, p. 11], it can further shown to be consistent with a thermal spectrum. To this end, assume that at any time, the NEH and the emitted radiation constitute a system in thermal equilibrium. Let $\lambda_{N,N-1}{\rm d}t$ denote the conditional probability that the system consists of the black hole plus $N$ quanta given that the system consisted of the black hole plus N1N-1 quanta a time ${\rm d}t$ ago. That is, $\lambda_{N,N-1}$ is the conditional probability per unit time of the system transitioning from the state ``black hole + (N1)(N-1) quanta'' to the state ``black hole + $N$ quanta'' by spontaneous emission. Similarly, $\lambda_{N-1,N}$ would be the conditional probability per unit time of the system transitioning from the state ``black hole + $N$ quanta'' to the state ``black hole + (N1)(N-1) quanta'' by absorption of a quantum. Now consider an ensemble of identical systems each defined as above, and let $\sigma_{N}$ be the probability that a randomly chosen system out of the ensemble is in the state ``black hole + $N$ quanta''. The principle of detailed balance implies that
$\frac{\lambda_{N,N-1}}{\lambda_{N-1,N}}=\frac{\sigma_{N}}{\sigma_{N-1}}=\exp\left(-8\pi M_{\text{ADM}}\leavevmode\nobreak\ \omega_{\infty}\right)\leavevmode\nobreak\ ,$ (V.3)
@@
$\braket{N}=\sum_{N=0}^{\infty}N\sigma_{N}=\frac{1}{e^{8\pi M_{\text{ADM}}\omega_{\infty}}-1}\leavevmode\nobreak\ ,$ (V.4)
which is a thermal spectrum of bosons at temperature $T=1/8\pi M_{\text{ADM}}$.
The above arguments generally cannot be straightforwardly generalized to non-singular gravitational collapse. While the outer horizon $H_{O}$ is non-expanding, and can thus be regarded as being in equilibrium, the inner horizon $H_{I}$, due to the bouncing geometry, necessarily evolves in a very nontrivial manner. Therefore, the trapping horizon may, on the whole, be far from equilibrium. Thus, only in the case ($\Delta=0$) of the ingoing antiparticle being absorbed by the black hole before reaching the inner horizon is the probability of pair-creation (IV.28) consistent with a thermal spectrum of radiation. For the case $\Delta=1$, there is a contribution to the probability from the inner horizon, which being far from equilibrium in general, makes the emitted radiation non-thermal as well. It might be possible to determine the precise deviation from thermality if there are regions of the inner horizon where the spacetime geometry is not changing very fast. A look at some explicit examples of non-singular collapse profiles [21, 30, 22] reveals that if the mass $M_{\text{ADM}}$ of the black hole is much larger than the Planck mass, the radius of the inner horizon evolves on the order of $\sim t$, except when it meets the outer horizon. Thus, at intermediary times, one may regard the inner horizon as approximately in equilibrium. In that case, one may invoke a local version of detailed balance and arrive at
$\braket{N}=\frac{1}{\exp\left(\frac{2\pi\omega_{H_{I}}}{\kappa_{H_{I}}}+\frac{2\pi\omega_{H_{O}}}{\kappa_{H_{O}}}\right)-1}\leavevmode\nobreak\ .$ (V.5)
Once again, we may write $\kappa_{H_{O}}=1/4M_{\text{ADM}}$ and also $\omega_{H_{O}}=\omega_{\infty}$ since the geometry outside the outer horizon is static. Thus,
$\braket{N}=\frac{1}{e^{8\pi M_{\text{ADM}}\omega_{\infty}}e^{2\pi\omega_{H_{I}}/\kappa_{H_{I}}}-1}\leavevmode\nobreak\ .$ (V.6)
Under the assumption of an inner horizon in approximate equilibrium, we may set $\dot{M}_{H_{I}}\approx 0$, whence $\kappa_{H_{I}}\approx-M^{\prime}_{H_{I}}/2M_{H_{I}}$ (cf. (II.14)). Thus, under our assumptions, the deviation from a thermal spectrum depends on how spatially inhomogeneous the inner horizon is. For very small inhomogeneities ($M^{\prime}_{H_{I}}\approx 0$), the spectrum of emitted radiation is almost thermal. However, this would be true of extremely fine-tuned mass profiles. In fact, a look at the complicated form of the mass function in Ref. [40] specially engineered to give a smooth bouncing profile suggests that it might not even be possible to construct a nearly homogeneous mass profile that yields the requisite bouncing geometry. In any case, it is certain that no mass profile can tranquilize the highly non-equilibrium behavior of the mass function in the vicinity of where the inner and outer parts of the trapping horizon meet. Therefore, in general, the non-singular collapse emits non-thermal radiation.
@@ VI Conclusions and discussion
Third, in certain LQG-inspired non-singular collapse scenarios based on a dust-gravity system, the shells of matter inevitably collide one another after the matter bounces [59]. The mass function $M(t,r)$ becomes discontinuous at such points since $\lim_{r\to S^{+}}M>\lim_{r\to S^{-}}$, where $r=S(t)$ denotes the location of the shell-crossing. Dynamics beyond a shell crossing have to be found out by integrating the equations of motion. This determines how the discontinuity in the mass function at the shell crossing evolves in time – such traveling discontinuities are called shock waves. There is, in general, an infinity of ways of integrating the equations of motion to find the evolution of the shock wave [75]. This is encapsulated in the so-called Rankine-Hugoniot jump condition for the speed of the shock wave,
$\frac{{\rm d}S}{{\rm d}t}=\frac{f(S,M^{+})-f(S,M^{-})}{g(S,M^{+})-g(S,M^{-})}\leavevmode\nobreak\ ,$ (VI.1)
where $M^{\pm}=\lim_{r\to S^{\pm}}M$ and $f,g$ are some functions of $r$ and $M$ whose precise form depends on the particular way that the equations of motion are integrated. In our description of the pair-creation process, we have ignored the possibility of shell crossings. It is worth inquiring about the implications of dropping this assumption. We offer some remarks to that end.
@@
Notice, first of all, that the fact that the matter bounces and re-expands in the same asymptotic region that it collapses in implies that the inner horizon must generally move in a spacelike manner. Therefore, since the ingoing antiparticle moves along a null geodesic, it must be absorbed by the matter in the rebound phase. Since, furthermore, shell crossings occur for generic initial data a Planck time after the bounce [59], the absorption must occur after the first instance of shell crossing. Finally, because the entire matter profile soon compresses into a thin shell of infinite energy density [30], the ingoing antiparticle is absorbed by the shock wave that forms after the bounce. Now, since the ingoing particle approaches the shock wave from the outside, the result of the absorption process must be a reduction in the value that the mass function function tends to from the right, i.e., $M^{+}$. Consequently, the jump in the mass function across the shock wave becomes smaller and this also changes the dynamics of the shock wave via the Rankine-Hugoniot condition. It is tempting to speculate that given enough number of ingoing quanta, the discontinuity should be entirely removed, smoothening out the mass function at the shock and thus dispensing with the shock wave altogether. This opens up the possibility of developing quantum-corrected models of non-singular gravitational collapse in a single asymptotic region without shell crossings. One avenue of exploring this possibility might be an LQG-corrected dust-gravity system as in Ref. [30] but coupled to a massless scalar field. Such an investigation is underway and the results will be reported in a future paper.
@@ Acknowledgments
This work owes its origin to numerous discussions with Viqar Husain. I am grateful to him, and also to Francesco Fazzini and Edward Wilson-Ewing for valuable feedback and discussions. This work was supported by the Natural Science and Engineering Research Council of Canada.

### v3

@@ IV Spontaneous emission as tunneling
    In this formalism, the propagator for a relativistic particle is given by $\braket{x|x^{\prime}}=\int d\lambda_{0}\penalty 10000\ e^{-im^{2}\lambda_{0}}F(\lambda_{0},x,x^{\prime})$, where $F(\lambda_{0},x,x^{\prime})$ is the amplitude of propagation from $x^{\prime}$ to $x$ in parameter time $\lambda_{0}$.
@@
    The paths for which the Hamilton-Jacobi equation holds are null. Tunneling involves traversing classically forbidden regions using complex paths (Refs. [57, Chapter VII] and [48]).
@@
Notice that in implementing step 5, we chose for the emitted particle a path which is close to the outer horizon (Fig. 2). This choice has a similar effect as the late-time approximation involved in the conventional calculation of Hawking radiation [39]. For in that calculation, the late-time approximation enables one to calculate the Bogoliubov coefficients between the late-time and early-time modes by solving the Klein-Gordon equation near the horizon. These coefficients in turn enter into the amplitude for there being particles at future null infinity at late times given that there were no particles incoming from past null infinity at early times. Somewhat similarly, the choice of a near-horizon path in the calculation above enables one to evaluate the residue (IV.26) and hence the imaginary part of the action, which then determines the probability of spontaneous emission. It is in this sense that the result (IV.28) is the quantum-corrected analog of the classical late-time result found by Hawking.
However, as can be seen in Fig. 2, in the quantum-corrected geometry, the outgoing particle whose path snuggles the outer horizon does not reach future null infinity at late times. There might well be particles emitted by the dynamical geometry that reach future null infinity after $u=u_{0}$, where $u_{0}$ is the retarded time at which the outgoing particle in Fig. 2 reaches future infinity. Now the spectra of these particles might have other corrections due to the highly nontrivial spacetime geometry to the future of $u_{0}$. But studying these corrections would require a knowledge of the precise details of the geometry in regions where the effective description of quantum-corrected collapse that we have employed may not be reliable (e.g., the bounce region). Therefore, our calculation sheds light only on any potential corrections to the spectrum of emitted particles which reach future infinity near $u_{0}$. These paths are close to the outer horizon, and as such enable us to abstract away the effect of the classical late-time approximation which is responsible for producing a thermal spectrum in the conventional case.
Before closing this section, a historical note is in order. The Hamilton-Jacobi method, encapsulated in (IV.3) and the picture of particle emission by a black hole as spontaneous pair creation, was first employed in the study of Hawking radiation by Srinivasan and Padmanabhan [70] and, independently, by Parikh and Wilczek [64], and has been considerably refined since then (see Ref. [78] for a comprehensive review). However, the typical starting point in all these works on black holes is the formula (IV.3), its origin left unexplained. In this and the preceding section, besides obtaining the probability of pair creation by non-singular gravitational collapse, an attempt has been made to provide the link between (IV.3) and the understanding of particle emission by black holes as spontaneous pair creation.
@@ V Comparison with the classical result
$\sigma\sim\exp\left(-\frac{2\pi\omega_{\text{NEH}}}{\hbar\kappa_{\text{NEH}}}\right)\penalty 10000\ .$ (V.1)
@@
$\sigma\sim\exp\left(-8\pi M_{\text{ADM}}\penalty 10000\ \omega_{\infty}/\hbar\right)$ (V.2)
By invoking the principle of detailed balance [78, p. 11], it can further shown to be consistent with a thermal spectrum. To this end, assume that at any time, the NEH and the emitted radiation constitute a system in thermal equilibrium. Let $\lambda_{N,N-1}{\rm d}t$ denote the conditional probability that the system consists of the black hole plus $N$ quanta given that the system consisted of the black hole plus N1N-1 quanta a time ${\rm d}t$ ago. That is, $\lambda_{N,N-1}$ is the conditional probability per unit time of the system transitioning from the state ``black hole + (N1)(N-1) quanta'' to the state ``black hole + $N$ quanta'' by spontaneous emission. Similarly, $\lambda_{N-1,N}$ would be the conditional probability per unit time of the system transitioning from the state ``black hole + $N$ quanta'' to the state ``black hole + (N1)(N-1) quanta'' by absorption of a quantum. Now consider an ensemble of identical systems each defined as above, and let $\sigma_{N}$ be the probability that a randomly chosen system out of the ensemble is in the state ``black hole + $N$ quanta''. The principle of detailed balance implies that
$\frac{\lambda_{N,N-1}}{\lambda_{N-1,N}}=\frac{\sigma_{N}}{\sigma_{N-1}}=\exp\left(-8\pi M_{\text{ADM}}\penalty 10000\ \omega_{\infty}/\hbar\right)\penalty 10000\ ,$ (V.3)
@@
$\braket{N}=\sum_{N=0}^{\infty}N\sigma_{N}=\frac{1}{e^{8\pi M_{\text{ADM}}\omega_{\infty}/\hbar}-1}\penalty 10000\ ,$ (V.4)
which is a thermal spectrum of bosons at temperature $T=\hbar/8\pi k_{B}M_{\text{ADM}}$.
The above arguments generally cannot be straightforwardly generalized to non-singular gravitational collapse. While the outer horizon $H_{O}$ is non-expanding, and can thus be regarded as being in equilibrium, the inner horizon $H_{I}$, due to the bouncing geometry, necessarily evolves in a very nontrivial manner. Therefore, the trapping horizon may, on the whole, be far from equilibrium. Thus, only in the case ($\Delta=0$) of the ingoing antiparticle being absorbed by the black hole before reaching the inner horizon is the probability of pair-creation (IV.28) consistent with a thermal spectrum of radiation. For the case $\Delta=1$, there is a contribution to the probability from the inner horizon, which being far from equilibrium in general, makes the emitted radiation non-thermal as well.
Under the assumption of an inner horizon in approximate equilibrium, we may set $\dot{M}_{H_{I}}\approx 0$, whence $\kappa_{H_{I}}\approx-M^{\prime}_{H_{I}}/2M_{H_{I}}$ (cf. (II.34)). Thus, under our assumptions, the deviation from a thermal spectrum depends on how spatially inhomogeneous the inner horizon is. For very small inhomogeneities ($M^{\prime}_{H_{I}}\approx 0$), the spectrum of emitted radiation is almost thermal. However, this would be true of extremely fine-tuned mass profiles. In fact, a look at the complicated form of the mass function in Ref. [47] specially engineered to give a smooth bouncing profile suggests that it might not even be possible to construct a nearly homogeneous mass profile that yields the requisite bouncing geometry. In any case, it is certain that no mass profile can tranquilize the highly non-equilibrium behavior of the mass function in the vicinity of where the inner and outer parts of the trapping horizon meet. Therefore, in general, the non-singular collapse emits non-thermal radiation.
@@ VI Conclusions and discussion
Third, in certain LQG-inspired non-singular collapse scenarios based on a dust-gravity system, the shells of matter inevitably collide one another after the matter bounces [26]. The mass function $M(t,r)$ becomes discontinuous at such points since $\lim_{r\to S^{+}}M>\lim_{r\to S^{-}}$, where $r=S(t)$ denotes the location of the shell-crossing. Dynamics beyond a shell crossing have to be found out by integrating the equations of motion. This determines how the discontinuity in the mass function at the shell crossing evolves in time – such traveling discontinuities are called shock waves. There is, in general, an infinity of ways of integrating the equations of motion to find the evolution of the shock wave [27]. This is encapsulated in the so-called Rankine-Hugoniot (RH) jump condition for the speed of the shock wave, an instance of which we saw in Section II. In our description of the pair-creation process, we have ignored the possibility of shell crossings. It is worth inquiring about the implications of dropping this assumption. We offer some remarks to that end.
@@
Notice, first of all, that the ingoing antiparticles which are created sufficiently long after the formation of the trapping region will be absorbed by the matter in the rebound phase. Under our assumption that the size of the trapping region $r_{H_{O}}\gg\sqrt{\hbar}$, there will be many such absorbed antiparticles. Since, furthermore, shell crossings occur for generic initial data a Planck time after the bounce [26], these antiparticles must be absorbed after the first instance of shell crossing. Finally, because the entire matter profile soon compresses into a thin shell of infinite energy density [51], the ingoing antiparticles are absorbed by the shock wave that forms after the bounce. Now, since an ingoing particle approaches the shock wave from the outside, the result of the absorption process must be a reduction in the value that the mass function function tends to from the right, i.e., $M^{+}$. Consequently, the jump in the mass function across the shock wave becomes smaller and this also changes the dynamics of the shock wave via the Rankine-Hugoniot condition. It is tempting to speculate that given enough number of ingoing quanta, the discontinuity should be entirely removed, smoothening out the mass function at the shock and thus dispensing with the shock wave altogether. This opens up the possibility of developing quantum-corrected models of non-singular gravitational collapse in a single asymptotic region without shell crossings.
Finally, it is worth re-emphasizing that the model of radiation from quantum-corrected collapse that has been presented above cannot be trusted in all the regions of spacetime. We do not have access to the precise details of the geometry of spacetime in a region of Planckian dimensions around the bounce. The effective field-theoretic studies on collapse may robustly establish that there is a bounce when matter reaches Planckian densities, but there is no guarantee that one can even describe gravity in terms of a smooth spacetime manifold at such scales. Therefore, all the calculations and conclusions in this paper apply to regions away from the quantum region where the bounce occurs. In particular, the probability of emission (IV.28) applies only to antiparticles which cross the inner horizon sufficiently long after the bounce (see Fig. 2). Keeping this in mind, in eqn.(IV.28), $\kappa_{H_{I}}$ and $\omega_{H_{{}_{I}}}$ are to be evaluated at times well after the bounce. For the quantum-corrected Oppenheimer-Snyder model in the framework of effective LQG, we saw in Section II that at such times, $\kappa_{H_{I}}$ can be approximated by eqn. (II.40). Plugging this into (IV.28), we see that the the inner-horizon correction to the conventional Hawking result looks like
$\frac{2\pi\omega_{H_{I}}}{\hbar(t-t_{0})}\penalty 10000\ .$ (VI.1)
This suggests that this correction is of the same order in $\hbar$ as the conventional term coming from the outer horizon. Therefore, to the extent that the semiclassical tunneling calculation can be trusted, the extra contribution from the inner horizon cannot be ignored as a higher-order effect which is inconsequential at the scales at which Hawking radiation is significant. Nevertheless, two shortcomings of the result derived in this paper must be clearly pointed. First, as we emphasized at the end of Section IV, the assumptions involved in our effective framework do not permit an analysis of spontaneous emission to the future of the outer horizon. Such emission might well have higher order corrections in its spectrum. Second, the result that we have derived can be fully trusted only if it can be reproduced from a more fundamental model of radiation from quantum-corrected collapse. That is, rather than having to put a scalar field on a putative model of collapse which involves quantum corrections to the geometry, such as a bounce instead of a central singularity, one should study a model that involves from the get-go any quantum corrections to both a matter-filled geometry and radiation from that matter. One avenue of developing such a model would be to add a scalar field to the dust-gravity system in the framework of effective LQG [51] before taking the semiclassical limit. At the effective level, this would result in coupled partial differential equations for the degrees of freedom of the scalar field and gravity. Solutions to these equations should incorporate corrections to both the geometry and the spectrum of the radiation due to the scalar field.
@@ Acknowledgments
This work owes its origin to numerous discussions with Viqar Husain. I am grateful to him, and also to Augusto Cabrera, Francesco Fazzini, Irfan Javed and Edward Wilson-Ewing for valuable feedback and discussions. This work was supported by the Natural Science and Engineering Research Council of Canada.